OpenStudy (anonymous):

Click on the symbol that will make this expression true. a = 4, b = 6, c = 8 5c - 3a - b + 12 ? 6 A. > B. < C.=

OpenStudy (anonymous):

40-12-6=22

OpenStudy (anonymous):

22+12=34 22+6=28

OpenStudy (anonymous):

34 is greater than 28

OpenStudy (anonymous):

so A.> would be the answer